# Fast, robust and laser-free universal entangling gates for trapped-ion quantum computing

### v1

@@ V Conclusion
In this work, we lay the theoretical and experimental framework for implementing RF driven double dressed state entangling gates with further increasing gate speed. Using this novel gate scheme, we create the Bell states $|\Phi^{+}\rangle$ and $|\Psi^{+}\rangle$ simply by choosing the appropriate interaction time of qubits with a phase-modulated driving field. Using this novel gate, fidelities up to $98^{+2}_{-3}\leavevmode\nobreak\ \%$ with gate time of $\leq 313\leavevmode\nobreak\ \mathrm{\mu}\text{s}$ were achieved, using readily available experimental parameters.
The qubits are double dressed by a single phase-modulated RF field which acts as continuous dynamical decoupling driving the entangling operation, and at the same time, protects the qubits’ coherence. Thus, this gate is intrinsically robust against RF field amplitude fluctuations, as well as magnetic field fluctuations that otherwise would adversely affect magnetically sensitive hyperfine-qubits.
@@ Appendix A Double dressed state of a single Qubit
We consider a single ion trapped in a harmonic potential. The Hamiltonian that describes two hyperfine states, given that the ion is subjected to a linear magnetic field gradient, is [7]
@@
We define the dressed basis operators by the canonical transformation $S_{x}=\sigma_{z},\ S_{y}=\sigma_{y},S_{z}=-\sigma_{x}$. The Hamiltonian in Eq. (13) then takes the form,
@@
$H_{qss}=-\frac{\left(\eta\nu\right)^{2}}{2\left(\nu-\Omega_{1}\right)}S_{z}b^{\dagger}b.$ (18)
@@
Appendix B Detection correction
The qubit readout is realized by collecting resonance fluorescence using a global laser beam close to $369$ nm. The fluorescence of each ion is imaged onto an EMCCD camera, producing a spatially resolved image of the ion chain. Using a two-threshold method to analyze the collected fluorescence, we distinguish the qubit’s logical states |$\ket{0}$ and |$\ket{1}$ [33]. For the experiments reported here, the product state probabilities $P_{i},\leavevmode\nobreak\ i=1,2,3,4$ of the computational basis states $\{1:\ket{00},2:\ket{01},3:\ket{10},4:\ket{11}\}$ need to be reconstructed. Due to imperfect spatial separation of the ions’ fluorescence, the readout may be assigned wrongly. To account for this detection error, we carry out a correction of the detected excitation probabilities of the product states. To this end, we prepare all possible computational basis states by first optically pumping into state |$\ket{00}$, then applying a Rapid Adiabatic Passage (RAP)[34] pulse, specific for the state to prepare, and finally read out the ions’ internal state. In this way, the probabilities of wrong assignments of readouts can be inferred. Here, $P_{1}$, $P_{2}$, $P_{3}$ and $P_{4}$ are the excitation probabilities of state |$\ket{00}$,|$\ket{01}$,|$\ket{10}$, and |$\ket{11}$ respectively. $\widetilde{P}$ is the measured excitation probability assigned using the double threshold detection. The state vector
$\vec{P}$ $=(P_{1},P_{2},P_{3},P_{4})\;,$ (19)
therefore, describes the excitation probability of all computational basis states. A linear map, $M$ between the real probability vector $\vec{P}$ and the apparent probability $\vec{\widetilde{P}}=M\vec{P}$ can be found. Here, $M$ is a real-valued $4\times 4$ matrix, and the elements of $M$ are given by the measured probabilities i.e $M_{ji}=\widetilde{P}_{i}$. This linear map, therefore, contains all possible wrong assignments. Applying the inverse matrix $\vec{P}=M^{-1}$ allows for reconstructing the excitation probabilities of the ions’ state, compensating for detection errors: $\vec{P}=M^{-1}\vec{\widetilde{P}}$. For the two-qubit system, we show as an example the corresponding detection matrix in Fig. 6. Assuming that all statistical errors follow a normal distribution, the error of the readout correction can be propagated following the descriptions in [35]. The standard deviation of the inverse matrix’s elements is
Appendix C Tomography
@@
$\rho$ $=\frac{1}{4}(\lambda^{(1)}_{1}\cdot\sigma_{z}\otimes\mathds{1}+\lambda^{(1)}_{2}\cdot\mathds{1}\otimes\sigma_{z}+\lambda^{(1)}_{3}\cdot\sigma_{z}\otimes\sigma_{z}$
@@ 3 $\pi$ $\langle\sigma_{y}\otimes\mathds{1}\rangle$ $\langle\sigma_{y}\otimes\sigma_{z}\rangle$
4 - $23\pi/$ 2 - $\langle\sigma_{x}\otimes\mathds{1}\rangle$ $\langle\sigma_{z}\otimes\sigma_{x}\rangle$
5 - $\pi$ $\langle\sigma_{y}\otimes\mathds{1}\rangle$ $\langle\sigma_{z}\otimes\sigma_{y}\rangle$
@@ Appendix D Negativity
Appendix E Gate evolution in the computational basis
@@
Appendix F Gate robustness
In a double dressed frame, addressing errors occur due to detunings $\delta\Omega_{1}$ of the dressed states energy gap and the second driving field’s frequency $\Omega_{1}$. In addition, the gate scheme requires a well-calibrated Rabi frequency $\Omega_{1}^{\text{Amp}}$, matching the chosen parameter $\Omega_{1}^{\text{Phase}}$, explained in Eq. (1). Experimentally, we investigate the robustness of the double dressed entangling gate against symmetric (asymmetric) detunings $\delta_{1}$ ($\delta_{2}$) of $\Omega^{\text{Amp}}_{1}$ w.r.t. $\nu-\epsilon$ by measuring the entangled state negativity at a gate time of 313 $\mathrm{\mu}$s with $\Omega_{1}^{\text{Phase}}=2\pi\times 94.8$ kHz. Fig. 9(a) shows the effect of a mismatch $\delta_{1}$ between the physical Rabi frequency $\Omega_{1}^{\text{Amp}}$ for both ions and the phase modulation frequency $\Omega_{1}^{\text{Phase}}$ ( $\delta_{1}=\Omega_{1}^{\text{Phase}}-\Omega_{1}^{\text{Amp}}$). Fig. 9(b) shows the impact of a mismatch of the physical Rabi frequencies between the two ions ($\delta_{2}=\Omega_{1}^{\text{Amp},1}-\Omega_{1}^{\text{Amp},2}$), while $\Omega_{1}^{\text{Amp},1}=\Omega_{1}^{\text{Phase}}$ holds. For $\delta_{1}\approx 3\leavevmode\nobreak\ \%$ and $\delta_{2}\approx 8\leavevmode\nobreak\ \%$ of $\Omega_{1}^{\text{Phase}}$, the measured negativity stays above $80\leavevmode\nobreak\ \%$ of the optimal negativity, showing the robustness of the entangling gate against such mismatches of the RF field amplitude. Similar results were observed for direct implemented MS gates in [38], quantified in terms of fidelity. For small detunings, as typically occur in experiments, a quadratic dependence between negativity and detuning is observed, showing the stability against small detunings. The robustness of the gate, made it possible to scan a wide range of $\delta_{1,2}$, exceeding typical uncertainties during the experiments by two orders of magnitude.

### v2

@@ V Conclusion
V.1 Phase modulation of RF drive
For implementing the two-qubit entangling gate described in this work, a single, phase-modulated RF field is applied to each qubit. In addition, we study the effect of this phase-modulated driving field on single qubit coherence (section III). Alternatively, for these purposes an amplitude-modulated field could be used [44, 57]. We discuss the advantages of phase modulation as compared to amplitude modulation for these purposes in what follows.
Cohen et al. [57] put forth two schemes for the generation of double-dressed states: (a) applying a second RF driving field that is on resonance with the dressed-state energy splitting determined by the first field driving the bare qubit states, or (b) by applying a second driving field that is detuned from the bare-state energy splitting by exactly the Rabi frequency. Here, we focus exclusively on case (b) due to its straightforward implementation with a single RF signal generated by an AWG.
In contrast, the realization of case (a) would necessitate the introduction of an additional external field with a frequency of $\Omega_{2}$ (close to the trap frequency $\nu$), which would then be responsible for creating the double-dressed states. It is important to note that this additional field would introduce additional amplitude and phase noise.
In case (b), two options are possible: b1) A phase-modulated implementation, which is used in this work, and b2) an amplitude-modulated implementation. In both of these approaches, the modulation of a single driving field is employed to generate sidebands around the carrier frequency instead of adding an independent driving field as in case (a). In this case the modulated signal, as guaranteed by the Jacobi-Anger expansion (e.g. [61]), ensures that the frequency components utilized to construct the gate are of equal amplitude relative to that of the carrier signal.
For phase modulation, the modulation depth $\Omega_{2}/\Omega_{1}$ can be as precise as the phase resolution of the signal applied to the qubit in the $\nu\approx 2\pi\times 12.6\leavevmode\nobreak$GHz regime. In this regime, a phase resolution of less than $\triangle\varphi=3.5\times 10^{-2}\leavevmode\nobreak$rad can be achieved through experimental means. Accordingly, at the modulation frequency $\Omega_{1}=2\pi\times 94.8\leavevmode\nobreak$kHz, a phase resolution of $\triangle\varphi=\triangle\varphi\cdot\Omega_{1}/\nu=3.5\times 10^{-7}\leavevmode\nobreak$rad can be achieved. As a consequence of the aforementioned resolution, a phase-modulated driving field permits the amplitude of the sidebands to be precisely calibrated with a relative step size of $10^{-5}$. Furthermore, the phase-modulated approach is immune against the non-linearity of RF amplifiers used for signal processing (see Appendix C). For these reasons, we believe that the phase-modulated implementation is a more promising avenue for future developments.
V.2 Summary and outlook
@@
In this work, we lay the theoretical and experimental framework for implementing RF driven double dressed state entangling gates with further increasing gate speed. Using this novel gate scheme, we create the Bell states $|\Phi^{+}\rangle$ and $|\Psi^{+}\rangle$ simply by choosing the appropriate interaction time of qubits with a phase-modulated driving field. We achieved fidelities up to $98^{+2}_{-3}\leavevmode\nobreak\ \%$ with gate time of $\leq 313\leavevmode\nobreak\ \mathrm{\mu}\text{s}$, using readily available experimental parameters. To our knowledge, this is the first experimental realization of an entangling gate based on double-dressed qubits.
The qubits are double dressed by a single phase-modulated RF field which acts as continuous dynamical decoupling driving the entangling operation, and at the same time, protects the qubits’ coherence. Thus, this gate is intrinsically robust against RF field amplitude fluctuations, as well as magnetic field fluctuations that otherwise would adversely affect magnetically sensitive hyperfine-qubits. Although the concrete implementation is platform-specific, double-dressing using phase modulation can be achieved in other physical systems.
@@ Appendix A Double dressed state of a single Qubit
In the main text, we give a brief summary of the theoretical concepts used to interpret the experimental results. Here, in the appendix, we give a more complete account of these. In order to make this appendix self-contained for the convenience of the reader, we repeat here expressions that are also used in the main text.
We consider driving a single ion co-trapped with N-1 ions in a 1-D harmonic potential. The Hamiltonian that describes two hyperfine states, given that the ion is subjected to a linear magnetic field gradient, is [9]
@@
The coupling between the internal and motional dynamics of the ion is determined by the effective Lamb-Dicke parameter $\eta$, defined as
$\eta=\frac{\triangle z\partial_{z}\omega(z)}{\nu}=g_{F}\mu_{B}(\partial_{z}B)\nu^{-3/2}/\sqrt{2Nm_{\text{Yb}}}$ (14)
Here $\triangle z$ is the extension of the motional ground state wave function. The qubit’s addressing frequency, $\omega(z)$, depends on the external magnetic field as described by the Breit-Rabi formula [64] for ${}^{171}\text{Yb}^{+}$
$E(z)$ $=E_{\text{HFS}}/4+g_{I}\mu_{N}B(z)m_{F}$ (15)
±$\pm E_{\text{HFS}}/2\left(1+2m_{F}\chi+\chi^{2}\right)^{1/2}$ (16)
$\chi$ $=\frac{(g_{J}+g_{I}m_{e}/m_{p})\mu_{B}B(z)}{E_{\text{HFS}}}=\frac{g_{F}\mu_{B}B(z)}{E_{\text{HFS}}}\;.$ (17)
Here, $E_{\text{HFS}}$ is the hyperfine splitting between levels with total angular momentum $F=I+1/2$ and $F=I-1/2$ in zero magnetic field, $\mu_{B}$ is Bohr’s magneton, $m_{e}$ and $m_{p}$ indicate the electron and proton mass, respectively, $g_{J}$ and $g_{I}$ are the electronic and nuclear g-factor, and the magnetic field is given by $B(z)=B_{0}+\delta_{z}B$. In equation 14 the linear Zeeman effect is considered.
@@
We define the dressed basis operators by the canonical transformation $S_{x}=-\sigma_{z},\ S_{y}=\sigma_{y},S_{z}=\sigma_{x}$. The Hamiltonian in Eq. (19) then takes the form,
@@
$H_{qss}=\frac{\left(\eta\nu\right)^{2}}{4\left(\nu-\Omega_{1}\right)}S_{z}b^{\dagger}b.$ (24)
@@
Appendix B Fidelity estimation for higher gradients
In the main text we claim that a higher magnetic field gradient will result in a faster gate with higher fidelity. In what follows, we estimate the different mechanisms that affect the fidelity in higher gradients. First, heating of the motional modes is considered. As was shown in the original work by Mølmer and Sørensen [48] this translates to an effective dephasing rate of
$1/T_{2}\approx\left(\frac{\eta\nu}{\nu-\Omega_{1}}\right)\frac{\gamma}{8},$ (25)
[…]
Note, that the coupling of the internal ion states to the motional mode in Ref. [48] is of the form $\sigma_{y}\left(b\text{e}^{-\text{i}\delta t}+b^{\dagger}\text{e}^{\text{i}\delta t}\right)$ while for the double-dressed states we have $F_{z}\left(b\text{e}^{-\text{i}\delta t}+b^{\dagger}\text{e}^{\text{i}\delta t}\right)$. Therefore, in our case the dephasing is along the Z axis, compared to the Y axis for a MS gate.
Increasing the gradient will also increase the effect of the off-resonance unitary terms and the unitary infidelity accordingly. The issue is the quantum Stark shift that affects the double-dressed states, as it leaves the ions entangled to the phonons at the end of the gate. In the following we drive an estimate for this shift.
Assuming, that the Rabi frequency $\epsilon\ll\nu$, we can approximate Eq. (22) to
$H_{II}\approx$ $\frac{\Omega_{2}}{4}F_{z}$ (26)
$-\frac{\eta\nu}{4}\left[F_{z}\left(b\text{e}^{-\text{i}\epsilon t}+b^{\dagger}\text{e}^{\text{i}\epsilon t}\right)-\text{i}F_{y}\left(b^{\dagger}\text{e}^{\text{i}\epsilon t}-b\text{e}^{-\text{i}\epsilon t}\right)\right].$
Moving to the rotating frame w.r.t. the double dressed state energy $\frac{\Omega_{2}}{4}F_{z}$, (26) transforms to
$H_{III}=$ $-\frac{\eta\nu}{4}F_{z}\left(b\text{e}^{-\text{i}\epsilon t}+b^{\dagger}\text{e}^{\text{i}\epsilon t}\right)$ (27)
$+\frac{\eta\nu}{4}\left(F_{+}\text{e}^{\text{i}\Omega_{2}t/2}-F_{-}\text{e}^{-\text{i}\Omega_{2}t/2}\right)\left(b^{\dagger}\text{e}^{\text{i}\epsilon t}-b\text{e}^{-\text{i}\epsilon t}\right).$ (28)
Note, that when $\Omega_{2}/2=\epsilon$, the flip-flop interaction is on-resonance. We suspect that this is the reason for the dip in the coherence times in Fig. 2. While the RWA $\epsilon\gg\eta\nu$ that justifies neglecting the rotating terms is not valid in the limit $\Omega_{2}\sim\epsilon=\eta\nu$, we still think it is helpful to understand the dynamics by the following argument. The dressed basis changes due to the Jaynes-Cummings interaction into a hybridized basis with the phonons. Then the expected coherence time can be approximated by the Jaynes-Cummings collapse time [67] $\sim 2\sqrt{2}/\eta\nu\sim 0.1\ \text{ms}$. Although deviations from this value are expected due to the crude nature of the approximation, we believe that the underlying reason for the dip in the coherence time is the same.
Returning to the operating regime of the gate, where $\left|\Omega_{2}/2\pm\epsilon\right|\gg\eta\nu/2$, the shift of the double-dressed state’s energy that depends on the phonon occupation can be estimated by the second-order effective Hamiltonian [65]
$H_{eff}$ $=\left(\frac{\eta\nu}{4}\right)^{2}\left[\frac{1}{\Omega_{2}/2-\epsilon}+\frac{1}{\Omega_{2}/2+\epsilon}\right]F_{z}b^{\dagger}b$ (29)
$\approx\frac{\left(\eta\nu\right)^{2}}{4\Omega_{2}}F_{z}b^{\dagger}b,$
where in the last transition we assumed that $\frac{\Omega_{2}}{2}\gg\epsilon=\nu-\Omega_{1}$. Denoting $\delta_{qss}=\frac{\left(\eta\nu\right)^{2}}{4\Omega_{2}}$, the gate dynamics can be understood by the following Hamiltonian
[…]
Assuming we start at the motional ground state, the first order Magnus expansion suggests that we create two distinct coherent states depending on the ions’ state:
$\alpha_{00}(t)=-\frac{\eta\nu}{2(\epsilon+2\delta_{qss})}\left[e^{i(\epsilon+2\delta_{qss})t}-1\right]$ (31)
$\alpha_{11}(t)=-\frac{\eta\nu}{2(\epsilon-2\delta_{qss})}\left[e^{i(\epsilon-2\delta_{qss})t}-1\right],$
while $\alpha_{01}=\alpha_{10}=0$. At the gate time we have
$\left|\alpha_{00}(t_{g})\right|^{2}=\left(\frac{\eta\nu}{\epsilon+2\delta_{qss}}\right)^{2}\sin^{2}\left(\delta_{qss}t_{g}\right)$ (32)
$\left|\alpha_{11}(t_{g})\right|^{2}=\left(\frac{\eta\nu}{\epsilon-2\delta_{qss}}\right)^{2}\sin^{2}\left(\delta_{qss}t_{g}\right)$
The infidelity can then be estimated as
$IF\approx$ $\frac{\left|\alpha_{00}(t_{g})\right|^{2}+\left|\alpha_{11}(t_{g})\right|^{2}}{4}=$ (33)
$\left[\left(\frac{\eta\nu}{\epsilon-2\delta_{qss}}\right)^{2}+\left(\frac{\eta\nu}{\epsilon+2\delta_{qss}}\right)^{2}\right]\frac{\sin^{2}\left(\delta_{qss}t_{g}\right)}{4}.$
Assuming that $\delta_{qss}\ll\epsilon$, we can simplify the expression to
$IF\approx\frac{\left(\delta_{qss}t_{g}\right)^{2}}{2}.$ (34)
For our current setup $\delta_{qss}\approx 0.03\leavevmode\nobreak\ \text{kHz}$, therefore, the expected infidelity is about $0.2\leavevmode\nobreak\ \%$. This estimation fits well with our simulations starting at the motional ground state. However, in our experiment we start with $\bar{n}\approx 0.6$ which results in a significant deviation from this value and the unitary infidelity reaches about $1.5\%$. Since this term increases approximately quadratically with the magnetic field gradient, it is expected to be the dominant source of error in the gate.
There are two ways to mitigate this error. The first is to work with the stretch mode instead of the COM mode. This will allow to increase $\Omega_{2}$ significantly (by a factor of $\sqrt{3}$ similar to the mode’s frequency $\nu_{\text{str}}=\sqrt{3}\nu_{\text{COM}}$) and therefore suppress this interaction efficiently. In particular, the Lamb-Dicke parameter (Eq. 14) decreases by $3^{-3/4}$, and the frequency $\delta_{qss}\sim(\eta\nu)^{2}/\Omega_{2}$ is reduced by $3^{-1}$. This scaling will lower the infidelity by $3^{-3/2}$ while the gate time only increases by a factor of $3^{1/4}$ due to $t_{g}\sim 1/(\eta\nu)$.
A second method for reducing this interaction is to employ additional pulse shaping techniques. Our simulation of the gate with the current experimental parameters and a standard Gaussian ramp of $\Omega_{1}$ demonstrates a notable reduction in unitary infidelity to approximately 0.1-0.2 %, while slightly increasing the gate duration to approximately 410-450 $\mathrm{\mathrm{\mu}s}$. It is expected that similar improvements can be achieved with higher gradients. More sophisticated pulse schemes are predicted to further enhance this reduction, potentially by an order of magnitude.
Appendix C RF amplitude calibration
The gate scheme requires a well-calibrated Rabi frequency $\Omega_{1}^{\text{Amp}}$, matching the chosen parameter $\Omega_{1}^{\text{Phase}}$, explained in Eq. (5). Experimentally, we calibrate the physical Rabi frequency $\Omega_{1}^{\text{Amp}}$ by well known Rabi measurements. We address each qubit by a not phase-modulated on-resonant RF field near $2\pi\times 12.6$ GHz. We measure the excitation probability for different interaction times, and fit the data using a sinusoidal function, thus extracting the Rabi frequency $\Omega_{1}^{\text{Amp}}$. By accordingly changing the RF amplitude, we calibrate $\Omega_{1}^{\text{Amp}}\approx\Omega_{1}^{\text{Phase}}$.
In order to efficiently carry out this calibration, we also measure the nonlinear dependence between the amplitude of the RF signal and the observed Rabi frequency. Addressing one qubit transition on resonance using the RF signal generated with an AWG, we measure Rabi oscillations with frequencies up to $2\pi\times 200$ kHz. We show the measured frequencies in Fig. 7. Saturation sets in for RF amplitudes at about half the maximum amplitude. For desired Rabi frequencies close to the trap frequency, we operate in this nonlinear regime which can also affects the usage of amplitude-modulated fields.
Figure 7: For one ion, we measure the on-resonant Rabi frequency $\Omega$ as a function of the amplitude $A$ of the RF signal. The measurement accuracy is better than $2\pi\times 0.1$ kHz. The non-linearity is due to saturation of the RF amplifiers.
Appendix D Tomography
@@
$\rho$ $=\frac{1}{4}(\mathds{1}\otimes\mathds{1}+\lambda^{(1)}_{1}\cdot\sigma_{z}\otimes\mathds{1}+\lambda^{(1)}_{2}\cdot\mathds{1}\otimes\sigma_{z}+\lambda^{(1)}_{3}\cdot\sigma_{z}\otimes\sigma_{z}$
@@ 3 $\pi$ $\langle\sigma_{y}\otimes\mathds{1}\rangle$ $\langle\sigma_{y}\otimes\sigma_{z}\rangle$
4 - $23\pi/$ 2 - $\langle\mathds{1}\otimes\sigma_{x}\rangle$ $\langle\sigma_{z}\otimes\sigma_{x}\rangle$
5 - $\pi$ $\langle\mathds{1}\otimes\sigma_{y}\rangle$ $\langle\sigma_{z}\otimes\sigma_{y}\rangle$
@@ Appendix E Negativity
Appendix F Gate evolution in the computational basis
@@
Appendix G Detection correction
The qubit readout is realized by collecting resonance fluorescence using a global laser beam close to $369$ nm. The fluorescence of each ion is imaged onto an EMCCD camera, producing a spatially resolved image of the ion chain. Using a two-threshold method to analyze the collected fluorescence, we distinguish the qubit’s logical states |$\ket{0}$ and |$\ket{1}$ [68]. For the experiments reported here, the product state probabilities $P_{i},\leavevmode\nobreak\ i=1,2,3,4$ of the computational basis states $\{1:\ket{00},2:\ket{01},3:\ket{10},4:\ket{11}\}$ need to be reconstructed from the measured relative frequencies $\widetilde{P_{i}}$ assigned using double threshold detection. Due to imperfect spatial separation of the ions’ fluorescence, the readout may be assigned wrongly. To account for this detection error, we carry out a correction of the detected excitation probabilities of the product states. To this end, we prepare all possible computational basis states by first optically pumping into state |$\ket{00}$, then applying a Rapid Adiabatic Passage (RAP)[69] pulse, specific for the state to prepare, and finally read out the ions’ internal state. In this way, the probabilities of wrong assignments of readouts can be inferred. The state vector
$\vec{P}$ $=(P_{1},P_{2},P_{3},P_{4})\;,$ (42)
therefore, describes the excitation probability of all computational basis states. A linear map, $M$ between the probability vector $\vec{P}$ and the measured relative frequencies $\vec{\widetilde{P}}=M\vec{P}$ can be found. Here, $M$ is a real-valued $4\times 4$ matrix, and the elements of $M$ are given by the measured relative frequencies, i.e $M_{ji}=\widetilde{P}_{i},\leavevmode\nobreak\ j=1,2,3,4$ corresponding to the prepared states. This linear map, therefore, contains all possible wrong assignments. Applying the inverse matrix $\vec{P}=M^{-1}$ allows for reconstructing the excitation probabilities of the ions’ state, compensating for detection errors: $\vec{P}=M^{-1}\vec{\widetilde{P}}$. For the two-qubit system, we show as an example the corresponding detection matrix in Fig. 10. Assuming that all statistical errors follow a normal distribution, the error of the readout correction can be propagated following the descriptions in [70]. The standard deviation of the inverse matrix elements is